Starting from rest, a 3.5 kg object is acted upon by an unbalanced force and acquires a speed of 14.0 m/s after moving a distance of 35 m. Find a.) the acceleration of the object; b.) the unbalanced force acting on the object, and c.) the time it takes for the object to attain a speed of 14. 0 m/s


1
Expert's answer
2020-11-25T06:59:04-0500

a)

"v_0=0,"

"s=\frac{v^2-v_0^2}{2a} \implies a=\frac{v^2}{2s}=\frac{14^2}{2\cdot35}=2.8" "\frac ms"

b)

"F=ma=3.5\cdot2.8=9.8 N"

c)

"v=v_0+at \implies t=\frac va=\frac{14}{2.8}=5 s"